# AstroPix: A Pixelated HVCMOS Sensor for Space-Based Gamma-Ray Measurement

### v1

@@ 4 Noise Studies
‘Noise’ hits occur when particles from background sources interact with the detector, such as cosmic ray interactions or naturally occurring background radiation. Fluctuations in the electronics can also trigger a comparator readout, even without a particle interaction. This is a ‘dark count’ and is measured as a rate, the dark count rate (DCR). The goal of these studies are to determine the percentage of pixels which are sensitive within the dynamic range, understand the dark count rate at different thresholds, and set an optimized threshold voltage such that DCR is minimized.
@@
The analog baseline, as read off the amplifier from selected pixels, shows random fluctuations up to mV over a nearly constant baseline. This constant noise floor is due to detector electronics and consists of shot noise from the sensor diode, and thermal and flicker noise from the electronics. The latter two sources are thought to be sub-dominant, as they carry a dependence on capacitance and the AstroPix_v3 pixel capacitance is roughly pF. Large fluctuations with a low rate are likely noise from the environment in which the sensor is operated. The dark count rate can be mitigated during data collection by setting the comparator threshold higher than mV above baseline. A keV photopeak from cadmium-109 is visible in measurements made in Sec. 5 with a mV threshold, so the mV fluctuations can be disregarded and do not impact the dynamic range. The impact of increasing threshold on DCR is shown in Fig. 7.
@@
Threshold values of mV still allow the detection of keV signals as required in Table 1 (see Section 5 for details). With nominal operational settings, % of pixels per array are too noisy for data collection and must be masked (their comparators disabled).
@@
The variation in dark count rate also illustrates a degree of variation between pixels within the same array. Process variations and device mismatch lead to this natural variation in amplifier gain, leading to pixels responses varying . Individual-pixel calibration therefore is required to correct for this variation. Section 5 outlines this calibration strategy and its results.
@@
The in-pixel comparator does not collect sub-threshold hits. Studies in Ref. 6 show that charge sharing between pixels does not induce above-threshold hits in neighboring pixels. However, the threshold is set to maximize the dynamic range and has been shown (in Sec. 5) to enable measurement below the keV requirement. In this way, some degree of charge splitting can be accounted for provided that there is enough to trigger a neighboring comparator operating at this low global threshold. Simulation studies are planned to investigate the potential impact of charge sharing and its effect on energy resolution.
@@
The studies of this section illustrate that DCR can be reduced to tolerable rates with AstroPix devices. This can be done by carefully setting a global threshold which minimizes DCR while still providing measurements within the required dynamic range of keV and identifying and masking noisy pixels with large DCRs. These noisy pixels have no geometric dependence within the array and have not shown induced charge sharing. With the optimal operation described here and standard run settings from Table 2, a single AstroPix_v3 array measures a total DCR below Hz. This measured rate is acceptable for future AstroPix applications including A-STEP and large-format future-observatories such as AMEGO-X. Additionally, future AstroPix versions are expected to decrease DCR with the ability to set thresholds at an individual pixel level [18].
@@ 5 Energy Resolution
A high bias voltage enables more complete depletion and in general more efficient charge collection. Here we consider an energy calibration conducted at V bias, as described in Ref. 16. The current draw off the bias line is nA888The IV curve from this study is shown in Fig. 5..
@@ 5.1 Energy Calibration and Resolution
The energy calibration procedure is described in detail in Ref. 16, and is applied to one Okmetic chip. After masking 8 noisy pixels, measurements of radioactive sources including Cd-109, Ba-133, Am-241, and Co-57 were made. Photopeaks could be identified from individual pixel spectra between keV. A feature indicative of a Compton edge at around keV was present in the data. Calibration curves for each pixel relate the expected photopeak energy and the mean measured ToT with a function of the form
where is the true photopeak energy and is the uncalibrated ToT value. The functional form involves a low-energy linear component and higher-energy exponential decay. This form is motivated by the in-pixel charge-sensitive amplifier which operates in two different gain regimes, creating a bilinear gain structure. The amplifier saturation effect is also reflected. One example calibrated pixel is shown in Fig. 8 and compared to an example pixel from AstroPix_v2. The difference in slope is due in part to differences in front-end amplifier tuning. Figure 9 shows a calibrated Am-241 spectrum for all 996 calibrated pixels after individual pixel calibration. The keV photopeak is clearly visible. Further work is underway to determine its source, though it is expected to be artificial.
When considering one array, of pixels achieve the low-energy floor requirement of keV sensitivity. of pixels meet the energy resolution requirement of keV at keV with a median full-width half-max (FWHM) of keV (). The lack of a confident feature identification at keV makes the measured dynamic range keV.
The high-end of the dynamic range is hindered in AstroPix_v3 due to incomplete depletion (see Section 5.2) and amplifier saturation. Simulations of the amplifier in “high dynamic range" mode showed that its output amplitude saturates at keV. This is consistent with observations which saw maximum energies (that could not be fully calibrated) of keV. A subsequent iteration of AstroPix implements test pixels with a dynamic feedback capacitance realized with an NFET device [20], which displays no saturation at energies over keV while maintaining performance through the full dynamic range. Additionally, more ToT bits are allotted to prevent a dynamic range loss due to overflow in the digitization step in future versions.
5.2 Impact of Depletion
The highest-resistivity AstroPix_v3 design (cm) was targeted for exploring the goal depletion depth of m within a V bias range. The two other resistivities are included for comparison with previous generations and consideration for further studies. The different depletion depths expected for each wafer resistivity is shown in Fig. 10.
@@ 5.2 Impact of Depletion
A capacitance-voltage (CV) curve of the cm Okmetic substrate is shown in Fig. 11. Notably, the function shape at low voltages is different from the linear dependence of on bias voltage. We attribute this to the depletion area development. At near-zero bias the depleted regions starts to grow from the implants, therefore the depleted area is smaller than the full chip size. For this reason, initially grows sub-linearly with the bias voltage until the depleted region between neighboring pixels reconnect. The other feature is a sharp capacitance change around V. It is dependent on the test frequency, indicating a possible relation to surface charge. Both features are under further investigation, however the functional shape of the CV curve suggests under-depletion and isolated depleted regions under each implant in the voltage range tested.
Estimations of the depletion depth that the Okmetic AstroPix_v3 can achieve was tested in Ref. 16 by measuring the count rates of X-ray sources. The method exploits a rather small path length of the daughter photo-electron in silicon, of less than m, leading to proportionality of the rate and the depleted volume. An average depleted depth is extracted using the pixel area. The drawback of this method is the need to rescale the rate to account for low-energy part of the photon spectrum. This indirect measurement results in m depletion at V bias and m depletion at V bias, which agrees with the PN junction model curve with the uncertainty in the resistivity. Fitting a PN junction model to the data reproduces a resistivity of cm.
A direct measurement of depletion depth from the same wafer using the Edge Transition Current Technique [22] was conducted at the Santa Cruz Institute for Particle Physics in November 2023. Current results from this study are detailed in Ref. 21 and Ref. 17. The results are consistent with Ref. 16 over the voltage range tested, however further analysis is ongoing.
@@ 6 Heavy-Ion Radiation Testing
A previous iteration of AstroPix, AstroPix_v2, was tested for radiation hardness and single-event effects (SEEs). The digital periphery design was not changed between the two versions. Like AstroPix_v3, the AstroPix_v2 chips tested were fabricated on an Okmetic (*cm) substrate. The presence of the on-chip digital periphery motivates this testing, as that area may be the most susceptible. Two classes of SEE were monitored:
@@
Single-event latchup (SEL), a catastrophic event leading to runaway power draws due to parasitic switching of the CMOS component transistors [23], and
@@
Single-event functional interrupt (SEFI), a temporary state where radiation causes bit flips which may degrade data or configuration but a system reset can restore nominal operation.
@@ 6.1 Experimental Setup
The radiation tolerance of AstroPix_v2 was tested in June 2022 at the Lawrence Berkeley National Laboratory (LBNL) Berkeley Accelerator Space Effects (BASE) Facility 88” cyclotron [24]. The beam provides a cocktail of ions with 16 MeV/amu tune shown in Table 3. A range of ions with differing atomic masses provides a range of testable linear energy transfer (LET), and tilting the detector plane relative to the ion beam direction of propagation provides an effective LET (the surface-incident LET in silicon at normal incidence divided by the cosine of the detector tilt angle). Effective range is the ion penetration depth in silicon calculated perpendicular to the surface of the sensor (penetration range multiplied by the cosine of the detector tilt angle).
Ion | Initial Energy | Air gap | Sensor tilt | Surface energy | Effective LET | Effective |
[MeV] | [mm] | angle [deg] | [MeV STD] | [MeVcm2/mg] | range [m] | |
40Ar+14 | 642 | 20 | 0 | 8.0 | 206 | |
63Cu+22 | 1007 | 20 | 0 | 19 | 141 | |
45 | 23 | 21 | 115 | |||
78Kr+28 | 1317 | 20 | 0 | 28 | 132 | |
45 | 23 | 32 | 107 | |||
124Xe+43 | 1975 | 20 | 0 | 57 | 98 | |
35 | 23 | 64 | 102 |
Spaceflight standards [25] require instruments to survive catastrophic latchup up to a LET of MeVcm2/mg and fluence of cm-2. Dosimetry is provided from BASE through calibration of a set of four photomultiplier tubes (PMTs) to a center PMT that is then removed during irradiation.
Figure 12 shows AstroPix_v2 on the left, mm away from the beam pipe. The chip carrier board and auxiliary boards are covered with aluminum to protect active electronic components from recoil ions. Digital data was collected from the full array (with of pixels masked to reduce noise), as well as analog data from the bottom left corner pixel. Four input voltage rails and the high voltage bias were also monitored during irradiation for SEL. Non-destructive SEFI events could present as features in either power draw or returned data.
High- and low-flux data sets were collected, where the average high flux /cm2/s and average low flux /cm2/s. The total fluence of high flux runs was /cm2 unless an destructive event occurred (see Sec. 6.2).
@@ 6.2 Radiation Results
No immediately destructive high-current events occurred, suggesting AstroPix_v2 has a destructive SEL threshold greater than MeVcm2/mg. The device is susceptible to recoverable soft errors with a LET threshold above MeVcm2/mg. It is therefore unlikely that proton-induced secondary ions will cause on-orbit SEEs, assuming a low-inclination Low-Earth Orbit.
Three potential classes of SEFI events were identified, and one ("Class B") has subsequently been identified as a software-based error in the DAQ code. The remaining two classes of events are potentially flipped bits in either the analog configuration ("Class A") or the returned digital data ("Class C") leading to loss of analog signal or digitized data corresponding to nonphysical pixel locations. Each class was corrected with a reset or power cycle of the chip.
Cross sections for each data collection run are calculated by fitting the data to a Weibull curve [26] of the form
@@
where is the effective LET, is the limiting or plateau cross section, is an onset parameter such that for , is a width parameter, and is a dimensionless exponent. Figure 13 shows conservative cross sections assuming all three event classes were SEFIs. Asterisks represent Class A events and blue filled circles represent Classes B and C. The light blue circles at high LET and low cross section are from runs with an average flux of /cm2/s, whereas the dark blue points were collected with an average flux of /cm2/s . The distinction is not made at lower LET, as this flux effect was not found with lower LET beams for similar fluences. To account for these flux effects when calculating an upper bound on-orbit SEFI rate, the data have been fit with two Weibull curves and event rates summed. Both curves consider all data with LET MeV*cm2/mg. Curve 1 additionally includes only dark blue (high flux) data above MeV*cm2/mg whereas curve 2 includes only light blue (low flux) high-LET data. The fits are intentionally worst-case with steep increases and saturation levels at the highest cross section data point as opposed to an average. Corresponding Weibull parameters are given in Table 4.
@@
1 | 18 | 5 | 1 | |
2 | 31 | 5 | 7 |
Using CREME96 [27], a worst-case galactic cosmic ray environment of interplanetary space at solar minimum was used to determine a SEFI rate. The sensitive volumes defined in CREME96 had x- and y-dimensions equal to the square root of the saturated cross section; each volume was assigned a depth of 2 m. Although the pixel depletion volume is approximately 70 m thick at the voltage utilized for testing, the CMOS circuitry that is potentially responsible for the SEFIs will have a shallower sensitive volume. Within CREME96, this shallower volume yields higher event rates than would a deeper sensitive volume designation, conservatively bounding the rate. Total event rates for the worst-case assumptions described above are given in Table 5.
@@
This result is unique to AstroPix_v2, though the digital block of AstroPix_v2 is identical to that of AstroPix_v3. The satisfactory results raise confidence in future designs. Future AstroPix versions will be tested in an identical manner to ensure continued compliance with spaceflight standards.

### v2

@@ 4 Noise Studies
‘Noise’ hits occur when particles from background sources interact with the detector, such as cosmic ray interactions or naturally occurring background radiation. Electronic fluctuations can trigger a comparator readout even without a particle interaction, and detector noise from leakage current can contribute to underlying noise levels. The goal of these studies is to quantify the impact of noise through electronics and detector noise sources by determining the percentage of pixels which are sensitive within the operational range, understand the noise rate at different thresholds, and setting an optimized threshold voltage such that the impact of noise is minimized.
The analog baseline, as read off the amplifier from selected pixels, shows random fluctuations up to mV over a nearly constant baseline. This constant noise floor is due to detector electronics and consists of shot noise from the sensor diode and thermal and flicker noise from the electronics. Shot noise is not dominant as AstroPix application is in low-radiation environments, though the chip’s digital periphery is robust against catastrophic radiation damage (see Sec. 6.2). Thermal and flicker noise are controlled due to the use of in-pixel low- and high-pass filtering respectively (Fig. 3). This filtering helps mitigate their dependence on pixel capacitance, as the pixel capacitance is roughly pF. Pixel capacitance is dominated by the capacitance of metal traces over each pixel and not by the diode capacitance. The trace routing will be optimized in further versions of AstroPix and is expected to decrease to fF per pixel. The HVCMOS nature of the AstroPix design necessarily carries higher capacitance values than standard CMOS pixels, however the optimized future design is aligned with other HVCMOS chips.
@@
Large fluctuations with a low rate are likely noise from the environment in which the sensor is operated. The dark count rate can be mitigated during data collection by setting the comparator threshold higher than mV above baseline. A keV photopeak from cadmium-109 is visible in measurements made in Sec. 5 with a mV threshold, so the mV fluctuations can be disregarded and do not impact the operational range. The impact of increasing threshold on noise is shown in Fig. 7.
@@
Threshold values of mV allow the detection of keV signals as required in Table 1 (see Section 5 for details). As shown in Sec. 5, a keV photopeak from cadmium-109 is visible with a mV threshold, so that the electronics fluctuations do not impact the operational range. With nominal operational settings (see Table 3), % of pixels per array are too noisy for data collection and must be masked (their comparators disabled).
@@
The variation in dark count rate also illustrates a degree of variation between pixels within the same array. Process variations and device mismatch lead to this natural variation in amplifier gain which cause some pixels to be sensitive to noise which remains below threshold in other pixels. This leads to pixel responses varying relative to one another555The variation is calculated by identifying a photopeak from a spectrum and fitting a Gaussian function to the histogram of all pixel responses to that photopeak. The reported pixel variation is the confidence interval associated with the fit. See Fig. 9.. Individual-pixel calibration therefore is required to correct for this variation. Section 5 outlines this calibration strategy and its results.
@@
The in-pixel comparator does not collect sub-threshold hits. AstroPix_v3 is designed to reduce charge sharing, or signal splitting between neighboring pixels due to ionizing particle positioning. Improperly accounting for charge sharing could reduce the measured energy resolution of the device, as there is unmeasured (sub-threshold) charge in the shared pixels. Studies in Ref. 5 show that any potential charge sharing between pixels does not induce above-threshold hits in neighboring pixels. However, the threshold is set to maximize the operational range and has been shown (in Sec. 5) to enable measurement below the keV requirement. In this way, some degree of charge splitting can be accounted for provided that there is enough to trigger a neighboring comparator operating at this low global threshold. Simulation studies are planned to investigate the potential impact of charge sharing and its effect on energy resolution.
@@
The studies of this section illustrate that noise can be reduced to tolerable rates with AstroPix devices. This can be done by carefully setting a global threshold which minimizes noise while still providing measurements within the required operational range of keV and identifying and masking noisy pixels with large noise rates. These noisy pixels have no geometric dependence within the array and have not shown induced charge sharing. With the optimal operation described here and standard run settings from Table 3, a single array measures a total noise rate below Hz. This measured rate is acceptable for future AstroPix applications including A-STEP and large-format future-observatories such as AMEGO-X.
@@ 5 Energy Resolution
A high bias voltage enables more complete depletion and in general more efficient charge collection. Here we consider an energy calibration conducted at V bias, as described in Ref. 16. The current drawn off the bias line is nA (Fig. 2).
@@ 5.1 Energy Calibration and Resolution
The energy calibration procedure is described in detail in Ref. 16, and is applied to one wafer 2 chip. After masking 8 noisy pixels, measurements of radioactive sources including Cd-109, Ba-133, Am-241, and Co-57 were made. Calibration curves for each pixel relate the expected photopeak energy and the mean measured ToT with a function of the form
where is the true photopeak energy and is the uncalibrated ToT value. The functional form involves a low-energy linear component and higher-energy exponential decay. This form is motivated by the in-pixel charge-sensitive amplifier which utilizes two feedback capacitances which manifest as different gain regimes, creating a bilinear gain structure.
Figure 8 shows the application of these calibration curves to spectra for all calibrated pixels after individual pixel calibration. The keV photopeak of Cd-109 is clearly visible, indicating that the necessary keV threshold is achieved.
Calibration precision is reflected in Fig. 9 where the full width half max (FWHM) of a given photopeak is shown for each calibrated pixel. This is illustrated as a function of photopeak energy in Fig. 10.
The median FWHM value for each photopeak is used to calculate energy resolution, defined as the ratio of the FWHM to the expected photopeak energy. This energy resolution dependence on energy can be found in Ref. 16.
When considering one array, of pixels achieve the low-energy floor requirement of keV sensitivity. of pixels meet the energy resolution requirement of keV at keV with a median full-width half-max (FWHM) of keV (). The lack of a confident Compton edge feature identification at keV makes the measured operational range keV.
The high-end of the operational range is hindered due to incomplete depletion (see Section 5.2) and amplifier saturation. Simulations of the amplifier in “high dynamic range” mode showed that its output amplitude saturates at keV. This is consistent with Fig. 8 where the Cs-137 spectrum is terminated around keV. A subsequent iteration of AstroPix implements test pixels with a dynamic feedback capacitance realized with an NFET device [21] to enable higher measurements before saturation. Additionally, more ToT bits are allotted to prevent an operational range loss due to overflow in the digitization step in future versions.
5.2 Depletion Depth Measurements
The highest-resistivity AstroPix_v3 design ( k in wafer 3) was targeted for exploring the goal depletion depth of m within a V bias range. The two other resistivities are included for comparison with previous generations and consideration for further studies. The different depletion depths expected for each wafer resistivity are shown in Fig. 11.
@@ 5.2 Depletion Depth Measurements
A capacitance-voltage (CV) curve of the wafer 2 substrate is shown in Fig. 12. Notably, the function shape at low voltages is different from the linear dependence of on bias voltage. We attribute this to the depletion area development. At near-zero bias the depleted regions starts to grow from the implants, therefore the depleted area is smaller than the full chip size. For this reason, we expect first to grow sub-linearly with the bias voltage until the depleted region between neighboring pixels reconnect. Indeed, the shape of the CV curve indicates a reduction of the slope below V with quasi-linear dependence above this voltage. The other feature is a sharp capacitance change around V. It is dependent on the test frequency, indicating a possible relation to surface charge. Both features are under further investigation, however the functional shape of the CV curve suggests under-depletion and isolated depleted regions under each implant in the voltage range tested.
Estimations of the depletion depth that wafer 2 chips can achieve was tested in Ref. 16 by measuring the count rates of X-ray sources. This indirect measurement results in m depletion at V bias and m depletion at V bias, which agrees with the PN junction model curve with the uncertainty in the resistivity. Fitting a PN junction model to the data reproduces a resistivity of .
A direct measurement of depletion depth from the same wafer using the Edge Transition Current Technique [23] was conducted at the Santa Cruz Institute for Particle Physics in November 2023. Current results from this study are detailed in Ref. 22 and Ref. 17. These preliminary studies estimate depletion depths of m at V which over-estimates the theoretically expected value from Fig. 11. Unaccounted for sources of error including reflections in the measurement setup, relative angle between the chip and the beam, and charge sharing could all account for the difference. Further analysis of these data is ongoing.
@@ 6 Heavy-Ion Radiation Testing
AstroPix has been targeted for use in -ray detectors in low-Earth orbits at low inclination or outside Earth’s radiation belts, requiring resilience to heavy-ion radiation environments. A previous iteration of AstroPix, AstroPix_v2, was tested for radiation hardness to single-event effects (SEEs). While minor changes in design or processing of CMOS semiconductor chips can have unintended consequences for SEE susceptibility, this experiment provides some indication for expected SEE response in future AstroPix revisions. Two classes of SEE were monitored:
@@
Catastrophic Single-event latchup (SEL), e . structive event leading to runaway power draws due to parasitic switching of the CMOS transistors [24], and
@@
Single-event upset (SEU), a temporary state where radiation causes bit flips which may degrade data or configuration. In the case of a single-event functional interrupt (SEFI), a system reset can restore nominal operation.
The design of the on-chip digital periphery, the most susceptible detector region, is identical in AstroPix_v2 and AstroPix_v3, and both chips were fabricated on a wafer 2 silicon substrate.
@@ 6.1 Experimental Setup
The radiation tolerance of AstroPix_v2 was tested in June 2022 at the Lawrence Berkeley National Laboratory (LBNL) Berkeley Accelerator Space Effects (BASE) Facility ” cyclotron [25]. The beam provides a cocktail of ions with MeV/amu tune shown in Table 4. Four ion species with differing atomic masses provide a range of linear energy transfers (LETs), and tilting the detector plane relative to the ion beam direction of propagation provided additional effective LETs (the surface-incident LET in silicon at normal incidence divided by the cosine of the detector tilt angle). Surface energy was calculated with the SRIM-2013 code [26]. The effective range reported is the penetration depth of the ion species in silicon (multiplied by the cosine of the detector angle from the ion beam perpendicular when tilting).
| Einitial | Air gap [mm] | Esurface | Eff. LET | Eff. | |
| Ion | range | ||||
| [MeV] | & Sensor tilt | [MeV] | [MeV cm2/mg] | [m] | |
| 40Ar+14 | 642 | 20 : 0∘ | 8.0 | 206 | |
| 63Cu+22 | 1007 | 20 : 0∘ | 19 | 141 | |
| 45 : 23∘ | 21 | 115 | |||
| 78Kr+28 | 1317 | 20 : 0∘ | 28 | 132 | |
| 45 : 23∘ | 32 | 107 | |||
| 124Xe+43 | 1975 | 20 : 0∘ | 57 | 98 | |
| 35 : 23∘ | 64 | 82 |
Figure 13 shows the experimental beamline set-up with AstroPix_v2 on the left, mm away from the beam pipe. The AstroPix chip carrier board and auxiliary boards are covered with aluminum to protect active electronic components from recoil ions. Digitized event data was collected from the full pixel array (with of pixels masked to reduce noise), as well as a diagnostic analog signal from a single pixel, and monitored in real-time for non-destructive SEEs. Four input voltage rails (analog V and V supplies, a V supply for the chip’s digital periphery, and a V supply for the auxiliary readout boards) and the high voltage bias were also recording during irradiation for destructive and non-destructive SEEs.
Spaceflight standards [27] require instruments to be robust against catastrophic latchup up to a LET between MeV*cm2/mg. Accurate flux measurements are provided from BASE through calibration of a set of four photomultiplier tubes (PMTs) to a centered PMT that is then removed before the beam run. Data were recorded for each beam exposure, beginning before the beam was on and ending shortly after the beam was shuttered. Each beam exposure was run to an cm-2 effective fluence, or until a SEFI event was recorded. For all effective LETs, a minimum of one beam exposure was performed to the full cm-2 effective fluence to test for destructive SELs, ignoring any SEFIs. Both high- ( cm-2s-1) and low-flux ( cm-2s-1) beam exposures were performed after it was noted that the higher beam exposure flux was overwhelming the digital readout.
@@ 6.2 Radiation Results
In total, data from beam exposures were collected across four ion species and two incident angles. Two chips were tested with one device used for exposures and the second used for the final three runs at the highest LETs for confirmation of results. During testing, three potential classes of non-destructive events were identified. The first type (“Class A”) is likely a SEFI caused by flipped bits in the on-chip configuration registers. Class A events were characterized by a loss of both digital data and diagnostic analog signal coupled with a drop in supplied V digital power to pre-configuration levels. Class A events were corrected with a digital chip reconfiguration without power cycling, and later replicated using a nm infrared laser to stimulate the digital periphery.
The other two types of non-destructive event (“Class B” and “Class C”) involved the processing of the detector’s digital data. Class B events were characterized by the off-chip digital decoder software failing to process unexpected data structure in the bitstream and crashing. Modifications to the digital decoder software during later analysis dramatically reduced the rate of decoder failures, but the underlying cause of the unexpected bitstream structure is unresolved. Event pileup due to high beam flux, issues with data encoding in the frontend firmware, or a true SEFI could all be potential culprits. A conservative approach has been taken to treat all Class B events as SEFIs for this analysis. Class C events are characterized by a bit flip that modified the leading byte of each digitized event data packet. This behavior would occasionally correct or restart itself during the run. Power cycling of the V analog input to the detector would correct the behavior. Class C event behavior was observed in pre-beam baseline noise data, but was not reproducible in later testing.
Table 5 catalogs the beam exposure runs by effective LET and the resulting SEFI class of the run if any occurred. No Class A SEFIs occurred below an LET of MeV*cm2/mg. Combined uncertainty arising from fluence and cross section uncertainties is estimated to s of beam flux. In the case of multiple runs at the same LET, the lowest fluence until SEE is recorded.
A cross section of cm2 indicates no observed SEFIs. It was observed that Class B events occurred much more frequently at low beam flux. At high flux it is believed that the detector digital output was likely overrun and not returning all event data.
| LET | Avg Beam Flux | Min Fluence | Cross Section | SEE Class | DUT |
| [MeVcm2/mg] | [cm-2s-1] | [cm-2] | [cm2] | ||
| 8 | None | 1 | |||
| 19 | None | 1 | |||
| 21 | A | 1 | |||
| 28 | B | 1 | |||
| 28 | A | 1 | |||
| 32 | C | 1 | |||
| 57 | B | 1 | |||
| 57 | A | 1 | |||
| 57 | B | 2 | |||
| 64 | B | 2 |
CREME96 [28], an engineering tool to calculate a worst-case galactic cosmic ray environment at solar minimum, was used to determine an on-orbit SEFI rate. For inputs into this tool, the SEFI cross-section as a function of LET was fit with a Weibull curve [29] of the form
where is the effective LET, is the limiting or plateau cross section, is an onset parameter such that for , is a width parameter, and is a dimensionless exponent. All three described classes of observed detector behavior were treated as SEFIs. The data, due to limited LET test points and the different LET-based behavior of the three event classes, does not lend itself to a Weibull fit. However, for the sake of a conservative parameter inputs for the CREME96 model, we apply an over-constrained fit assuming worse possible factors to conservatively bound the on-orbit rate.
@@
To account for these worst-case flux effects when calculating an upper bound on-orbit SEFI rate, the data have been fit with two Weibull curves and event rates summed. One curve was fit to all data below LET MeV*cm2/mg and data from the high-flux runs above LET MeV*cm2/mg. The second curve was added to include the low flux results at high LET that have high cross section, and using a conservative onset LET of MeV*cm2/mg, as SEE cross section established at LETs of , , and MeV*cm2/mg remained level. The fits are intentionally worst-case with steep increases and saturation levels at the highest cross section data point as opposed to an average. Corresponding Weibull parameters are given in Table 6.
@@
| 1 | 18 | 5 | 1 | |
| 2 | 31 | 5 | 7 |
The sensitive volumes defined in CREME96 had x- and y-dimensions equal to the square root of the saturated cross section; each volume was assigned a depth of m. The CMOS circuitry that is potentially responsible for the SEFIs have a shallow sensitive volume relative to chip thickness. Within CREME96, this shallower volume yields higher event rates than would a deeper sensitive volume designation, conservatively bounding the rate. Total event rates for the worst-case assumptions described above are given in Table 7.
@@
After completing beam exposures at all effective LETs, no destructive high-current events occurred, suggesting AstroPix_v2 has a destructive SEL threshold greater than MeV*cm2/mg. The device is susceptible to recoverable soft errors with a LET threshold above MeV*cm2/mg. The onset LET above MeV*cm2/mg indicates proton-induced secondary ions are unlikely to cause on-orbit SEEs. This result is unique to AstroPix_v2, though the digital block of AstroPix_v2 is identical to that of AstroPix_v3. The satisfactory results raise confidence in future designs. Future AstroPix versions will be tested in a similar manner to ensure continued compliance with spaceflight standards.